NEW POTENTIAL ENERGY SURFACES FOR THE X̃ AND Ã STATES OF CH_{2}^{+}

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; W. P. KRAEMER, Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Strasse 1, Postfach 1523, D–85740 Garching, Germany; S. N. YURCHENKO^{a} and W. THIEL, Max-Planck-Institut für Kohlenforschung, Kaiser-Wilhelm-Platz 1, D–45470 Mülheim an der Ruhr, Germany; C. F. NEESE and J. L. GOTTFRIED, Department of Chemistry, The University of Chicago, Chicago, Illinois 60637; and PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

We report new ab initio calculations of the three-dimensional potential energy surfaces for the Renner-effect coupled X̃^{2}A_{1} ground electronic state and Ã^{2}B_{1} first excited electronic state of the CH_{2}^{+} molecule. We also make an ab initio calculation of the spin-orbit coupling surface A_{SO}(r_{12}, r_{32}, ρ) between these states. Using these ab initio surfaces, and our previously obtained^{b} ab initio dipole moment and transition moment surfaces, in our computer program RENNER, we calculate term values and absorption line intensities. We compare with recently observed high resolution spectra. Adjusting two parameters in the potential surfaces we are able to achieve satisfactory agreement with the experimental results except for those that involve the Ã state (v_{2}^{linear} = 8, l = 1) vibronic level. The implication of this disagreement is discussed. Apart from having the facility of using dipole moment and transition moment surfaces so that absolute intensity calculations can be made, our RENNER computer program has the second unique feature of being able to calculate the energies of high lying rotational states because it is based on the Hougen-Bunker-Johns Hamiltonian. This has made it possible to study the phenomenon of rotational energy level clustering in Renner molecules. A recent study for PH_{2} exemplifies this aspect of our work^{c}

aPresent address: Institut fur Physikalische Chemie und Elektrochemie, TU Dresden, D-01062 Dresden, Germany

bG. Osmann, P. R. Bunker, P. Jensen, and W. P. Kraemer, Chem. Phys. 225, 33-54 (1997).

cS. N. Yurchenko, W. Thiel, P. Jensen, and P. R. Bunker, J. Mol. Spectrosc. 239, 160-173 (2006).